# Gromov-Wasserstein Barycenters: The Analysis Problem

### v1

@@ 1. Introduction
These models allow for synthesis and analysis of points in a metric space with distance $dist(\cdot,\cdot)$, through variational problems relative to given reference points $\mathbf{x}^{1},\dots,\mathbf{x}^{S}$. For the synthesis problem, given a vector of non-negative coefficients $\lambda=(\lambda_{1},\dots,\lambda_{S})$ in the $(S-1)$-dimensional simplex $\Delta_{S-1}$, one seeks a point in the space minimizing the following objective:
$\mathbf{x}_{\lambda}\in\operatorname*{arg\,min}_{\mathbf{x}}\sum_{s=1}^{S}\lambda_{s}\,dist^{2}(\mathbf{x}^{s},\mathbf{x});$ (1)
@@
$\lambda_{\mathbf{y}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}^{2}(\mathbf{y},\mathbf{x}_{\lambda}),$ (2)
where $d_{a}(\cdot,\cdot)$ is a discrepancy function (e.g., $d_{a}=dist$ or another metric in the space or even a weaker notion of dissimilarity that still identifies indiscernibles). We refer to the collection of points generated via (1) for fixed templates $\{\mathbf{x}^{s}\}_{s=1}^{S}$ as a barycentric coding model (BCM). Consequently, the barycenter of a set of templates for given weights $\lambda$ provides an alternative way of combining elements, one that better reflects the underlying geometry of the space while still preserving the dimensionality reduction properties of the representation.
This framework has been extensively used in classical optimal transport (OT), where the space of probability measures is endowed with the Wasserstein distance [49, 64, 65]. Foundational works on Wasserstein barycenters include [1, 18, 8, 32, 71], with more recent developments such as [68, 25, 35, 69] and extensions to dictionary learning [51, 41].
As an illustration, consider the space of discrete probability measures with finite support on $\mathbb{R}$, equipped with the 2-Wasserstein metric $W_{2}(\cdot,\cdot)$. Let the template measures be $\delta_{0}$ and $\delta_{1}$, where $\delta_{x}$ denotes the Dirac measure centered at $x$. The corresponding barycenter space consists of the set $\{\delta_{t}:\,t\in[0,1]\}$, which is obtained by solving the synthesis problem (1) in this specific setting. In other words, each measure $\mu=\delta_{t}$ minimizes the functional $\mu\mapsto t\,W_{2}^{2}(\delta_{0},\mu)+(1-t)\,W_{2}^{2}(\delta_{1},\mu)$. Hence, we can interpret the coordinates of the measure $\delta_{t}$ with respect to the templates $\delta_{0}$ and $\delta_{1}$ as the vector $(t,1-t)$ in $\Delta_{1}$. In this simple example, barycenters correspond to convex combinations of the points 00 and $1$–that is, Euclidean barycenters. In more general settings, however, measures or distributions have the ability to represent more complex data than individual points in space. For instance, Figure 1 compares interpolation between two weighted point clouds using Wasserstein geometry versus Euclidean geometry. Interpolating between two templates can be understood as constructing the barycentric space generated by those reference templates. Specifically, if $\mu_{0}$ and $\mu_{1}$ are the probability distributions corresponding to two templates, and $\pi$ is an OT plan between them in the 2-Wasserstein sense, then the interpolated measure $\mu_{t}$ defined as the pushforward of $\pi$ under the map $(x,y)\mapsto(1-t)x+ty$, that is, $\mu_{t}:=((1-t)x+ty)_{\#}\pi$ for $t\in[0,1]$, coincides with the Wasserstein barycenter of $\mu_{0}$ and $\mu_{1}$ with respective weights $(1-t)$ and $t$:
$\mu_{t}=\operatorname*{arg\,min}_{\mu}\left\{(1-t)W_{2}^{2}(\mu,\mu_{0})+tW_{2}^{2}(\mu,\mu_{1})\right\}.$ (3)
Figure 1. Illustration of different barycenter spaces between two point clouds selected from the Point Cloud MNIST 2D dataset [23], shown for various $t\in[0,1]$ with interpolation coordinates $(1-t,t)$ on the horizontal axis. Top: OT interpolations (with respect to the 2-Wasserstein geometry). Bottom: Linear interpolations (with respect to the Euclidean geometry).
@@
These innovations have led to applications across diverse domains, ranging from quantum chemistry [46] to natural language processing [3], enhancing fundamental tasks such as clustering [15], dimensionality reduction [61], shape correspondence [30], shape analysis [39], computer vision [52], and inference-based simulation [28]. Extensions to unbalanced settings have been explored [5, 9, 34, 54, 75], along with other variants, such as Sliced GW [58], and Fused GW–a simultaneous extension of Wasserstein and GW distances–[63, 62, 57]. However, GW distance computation is challenging due to requiring to solve a non-convex quadratic program; approximations include linearized [7] and entropy-regularized [46, 55, 48] GW.
@@
$\Delta_{S-1}\ni\lambda\longmapsto\mathbb{Y}_{\lambda}\in\operatorname*{arg\,min}_{\mathbb{Y}}\sum_{s=1}^{S}\lambda_{s}\,GW^{2}(\mathbb{X}^{s},\mathbb{Y}).$ (4)
We call the collection generated by (4), for fixed templates and varying coefficients $\lambda\in\Delta_{S-1}$, a GW barycenter space or Gromov–Wasserstein barycentric coding model (GW-BCM).
One of the main differences between the 2-Wasserstein distance and the GW distance lies in how they support barycenter constructions. In the $W_{2}$ setting, barycenters between two or more measures can be constructed via displacement interpolation using OT maps under regularity assumptions [36, 1]. This interpolation aligns mass in a common ambient space and is underpinned by the pseudo-Riemannian structure of Wasserstein space [71]. In contrast, GW distance compares measures on different metric spaces by optimizing over structural correspondences rather than spatial displacements. While geodesics between two metric measure spaces exist in GW space and can yield meaningful barycenters (see Figure 2), there is no analogous displacement interpolation for more than two templates. This reflects the absence of a pseudo-Riemannian framework (in the classical sense [42, 4]), limiting the direct geometric interpretation of GW barycenters.
Figure 2. Illustration of GW barycenters between two point clouds from the Point Cloud MNIST 2D dataset [23], shown for various $t\in[0,1]$ with interpolation coordinates $(1-t,t)$ on the horizontal axis. The top subplot uses the function ot.gromov.gromov_barycenters from the POT Library [21] (based on [46]). The bottom subplot uses the blow-up technique from [13], which appropriately realigns the nodes and enlarges the template matrices $\mathbf{X}^{0}$ (zero shape) and $\mathbf{X}^{1}$ (one shape), creating new versions $\mathbf{X}^{0}_{b}$ and $\mathbf{X}^{1}_{b}$ of the same size, so that the GW barycenters can be interpreted as convex combinations $(1-t)\mathbf{X}^{0}_{b}+t\mathbf{X}^{1}_{b}$ (see Remark B.9 in App. B.3). MDS embedding is used for visualization.
In this article, we focus on estimating the unknown GW barycentric coordinates of a point (or, more precisely, a network) under the barycentric coding model within the Gromov–Wasserstein framework (GW-BCM), a task we refer to as the GW analysis problem. That is, given a network $\mathbb{Y}$ we aim to formalize and solve
$\lambda_{\mathbb{Y}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}^{2}(\mathbb{Y},\mathbb{Y}_{\lambda}),$ (5)
where $\mathbb{Y}_{\lambda}$ is generated by the GW-BCM (4), for some appropriate dissimilarity metric $d_{a}(\cdot,\cdot)$. In other words, assuming a given network is a GW barycenter of a set of templates $\{\mathbb{X}^{s}\}_{s=1}^{S}$, our goal is to recover the corresponding vector of weights $\lambda$. See Figure 2 for an illustration.
@@ 1.1. Summary of Contributions
    A gradient-based method motivated by an interpretation of GW barycenters not necessarily as exact minimizers of (4), but rather as critical points of a functional on the GW space.
Regarding (1), we adopt the following simple idea, which can be applied beyond the context of GW. Consider a finite set of templates $\mathbb{X}^{s}=(\mathbf{X}^{s},\mathrm{p}^{s})$, $s=1,\dots,S$. Given a weight vector $\lambda\in\Delta_{S-1}$, consider GW synthesis problem (4), where the minimization is performed over $\mathbb{Y}=(\mathbf{Y},\mathrm{q})$, with fixed matrix size $M$ for $\mathbf{Y}$ and a fixed probability vector $\mathrm{q}$ [46]. Assume that the GW synthesis problem (4) can be solved via a fixed-point iteration of the form $\mathbf{Y}_{n}=\rho_{\lambda}(\mathbf{Y}_{n-1})$. Then, if $(\mathbf{Y},\mathrm{q})$ lies in the GW barycenter space generated by the templates, the matrix $\mathbf{Y}$ must be a fixed point of the corresponding iterative scheme. That is, there exists an unknown ‘coordinate’ vector $\lambda_{\mathbf{Y}}$ such that $\mathbf{Y}=\rho_{\lambda_{\mathbf{Y}}}(\mathbf{Y})$. To recover $\lambda_{\mathbf{Y}}$, we differentiate a chosen matrix norm $\|\mathbf{Y}-\rho_{\lambda}(\mathbf{Y})\|$ with respect to $\lambda$ and find its minimum at $\lambda=\lambda_{\mathbf{Y}}$. In summary, to generate GW barycenters, we employ a simplified version of the iterative scheme proposed in [46], and show in Theorem 3.3 that elements of the barycenter space are fixed points of this iteration. Building on this, Theorem 3.6 provides an exact recovery method for $\lambda_{\mathbf{Y}}$, for any $\mathbf{Y}$ in the GW barycenter space, formulated as a convex quadratic program and implemented in Algorithm 2. See Figure 3 for an illustration of its applicability.
@@ 1.2. Organization of the Paper
Section 2 reviews the GW distance. Our main results, their proofs, and new algorithms are presented in Section 3 and Section 4, which address the GW analysis problem via two approaches: a fixed-point method (3) and a gradient-based method (4). 5 presents experiments on 2D and 3D point clouds that serve as a proof-of-concept for the analytical results. Conclusions follow in 6. The Appendix includes: additional context on the BCM framework in App. A; background proofs and examples in App. B; computational complexity of the proposed GW analysis algorithms in App. C; further experimental results in App. D; and an additional discussion on the computation of GW barycenters and $W_{2}$-barycenters in App. E. The code to reproduce the experiments of this paper is available online at https://github.com/RocioDM/GW_BCM.
@@ 1.3. Notation
In general, superscripts will be used for the templates and subscripts for the synthesized objects. We denote by $[S]:=\{1,\dots,S\}$ the set of indices from $1$ to $S$. The space of weights $\Delta_{S-1}$, also known as a standard simplex or probability simplex, is
$\Delta_{S-1}:=\{\lambda=(\lambda_{1},\dots,\lambda_{S})\in\mathbb{R}^{S}\mid\lambda_{s}\geq 0\,\,\,\,\forall s\in[S],\text{ and }\sum_{s\in[S]}\lambda_{s}=1\}.$
If $\mathrm{p}$ is a vector and $\mathbf{X}$ a matrix, when there is no ambiguity regarding whether we are referring to entries or coordinates, we will write $\mathrm{p}[i]=\mathrm{p}_{i}$ and $\mathbf{X}[i,j]=\mathbf{X}_{ij}$. We recall that the Frobenius norm of a real-valued square matrix $\mathbf{X}$ is defined as $\|\mathbf{X}\|_{\mathrm{Frob}}:=\sqrt{\mathrm{tr}(\mathbf{X}^{T}\mathbf{X})}$, where $\mathbf{X}^{T}$ stands for transpose. The component-wise (Hadamard) product between matrices is denoted by $(\mathbf{X}\odot\mathbf{Y})_{jl}:=\mathbf{X}_{jl}\mathbf{Y}_{jl}.$ For a vector $\mathrm{q}$, the matrix $\frac{1}{\mathrm{q}\mathrm{q}^{T}}$ is defined by $\left(\frac{1}{\mathrm{q}\mathrm{q}^{T}}\right)[{j,l}]:=\frac{1}{\mathrm{q}_{j}\mathrm{q}_{l}}.$
Given two (Borel) measure spaces $(X,\mu_{X})$ and $(Y,\mu_{Y})$, we say that the measure $\mu_{Y}$ is the pushforward of the measure $\mu_{X}$ by a measurable map $\varphi:X\to Y$, denoted by $\varphi_{\#}\mu_{X}=\mu_{Y}$, if $\mu_{Y}(A)=\mu_{X}(\{x\in X\mid\varphi(x)\in A\})$ for every measurable set $A\subseteq Y$. Given $\omega:Y\times Y\to\mathbb{R}$, we use the pullback notation $\varphi^{*}\omega(x,x^{\prime}):=\omega(\varphi(x),\varphi(x^{\prime}))$, for all $x,x^{\prime}\in X$. A coupling or transport plan $\pi$ between $\mu_{X}$ and $\mu_{Y}$, denoted $\pi\in\Pi(\mu,\nu)$, is a probability measure on the product space $X\times Y$ with marginals $\mu_{X}$, $\mu_{Y}$. In the case of finite discrete measures, $\mu_{X}=\sum_{i=1}^{N}\mathrm{p}_{i}\delta_{x_{i}}$ and $\mu_{Y}=\sum_{j=1}^{M}\mathrm{q}_{j}\delta_{y_{j}}$, each coupling can be identified with a matrix in
$\Pi(\mathrm{p},\mathrm{q}):=\{\pi\in\mathbb{R}^{N\times M}\mid\,\pi_{ij}\geq 0,\,\forall i,j\,;\,\sum_{j\in[M]}{\pi_{ij}}=\mathrm{p}_{i}\,\forall i\,\text{ and }\sum_{i\in[N]}\pi_{ij}=\mathrm{q}_{j}\,\forall j\}.$

### v2

@@ 1. Introduction
These models allow for synthesis and analysis of points in a metric space with distance $dist(\cdot,\cdot)$, through variational problems relative to given reference points $\mathbf{x}^{1},\dots,\mathbf{x}^{S}$. For the synthesis problem, given a vector of non-negative coefficients $\lambda=(\lambda_{1},\dots,\lambda_{S})$ in the (S1)(S-1)-dimensional simplex $\Delta_{S-1}$, one seeks a point in the space minimizing the following objective:
(1) $\mathbf{x}_{\lambda}\in\operatorname*{arg\,min}_{\mathbf{x}}\sum_{s=1}^{S}\lambda_{s}\,dist^{2}(\mathbf{x}^{s},\mathbf{x});$
@@
$\lambda_{\mathbf{y}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}^{2}(\mathbf{y},\mathbf{x}_{\lambda}),$
where $d_{a}(\cdot,\cdot)$ is a discrepancy function (e.g., $d_{a}=dist$ or another metric in the space or even a weaker notion of dissimilarity that still identifies indiscernibles). We refer to the collection of points generated via (1) for fixed templates $\{\mathbf{x}^{s}\}_{s=1}^{S}$ as a barycentric coding model (BCM). Consequently, the barycenter of a set of templates for given weights $\lambda$ provides an alternative way of combining elements, one that better reflects the underlying geometry of the space while still preserving the dimensionality reduction properties of the representation. This framework has been extensively used in classical optimal transport (OT), where the space of probability measures is endowed with the Wasserstein distance [52, 67, 68]. See Section A.1 in Suppl. Mat. Foundational works on Wasserstein barycenters include [1, 19, 8, 33, 77], with more recent developments such as [72, 26, 37, 73] and extensions to dictionary learning [54, 44].
@@
These innovations have led to applications across diverse domains, ranging from quantum chemistry [49] to natural language processing [3], enhancing fundamental tasks such as clustering [16], dimensionality reduction [64], shape correspondence [31], shape analysis [40], computer vision [55], and inference-based simulation [29]. Extensions to unbalanced settings have been explored [5, 10, 36, 57, 81], along with other variants, such as Sliced GW [61], and Fused GW–a simultaneous extension of Wasserstein and GW distances–[65, 66, 60]. However, GW distance computation is challenging because it requires solving a non-convex quadratic program; approximations include linearized [6] and entropy-regularized [49, 58, 51] GW.
@@
(2) $\Delta_{S-1}\ni\lambda\longmapsto\mathbb{Y}_{\lambda}\in\operatorname*{arg\,min}_{\mathbb{Y}}\sum_{s=1}^{S}\lambda_{s}\,GW^{2}(\mathbb{X}^{s},\mathbb{Y}).$
We call the collection generated by (2), for fixed templates and varying coefficients $\lambda\in\Delta_{S-1}$, a GW barycenter space or Gromov-Wasserstein barycentric coding model (GW-BCM).
In this article, we focus on estimating the unknown GW barycentric coordinates of a point (or, more precisely, a network) under the barycentric coding model within the Gromov-Wasserstein framework (GW-BCM), a task we refer to as the GW analysis problem. That is, given a network $\mathbb{Y}$ we aim to formalize and solve
(3) $\lambda_{\mathbb{Y}}\in\operatorname*{arg\,min}_{\lambda\in\Delta_{S-1}}d_{a}^{2}(\mathbb{Y},\mathbb{Y}_{\lambda}),$
where $\mathbb{Y}_{\lambda}$ is generated by the GW-BCM (2), for some appropriate dissimilarity metric $d_{a}(\cdot,\cdot)$. In other words, assuming a given network is a GW barycenter of a set of templates $\{\mathbb{X}^{s}\}_{s=1}^{S}$, our goal is to recover the corresponding vector of weights $\lambda$. See Fig. 1 for an illustration.
Figure 1. Illustration of GW barycenters between two point clouds from the Point Cloud MNIST 2D dataset [24], shown for various $t\in[0,1]$ with interpolation coordinates (1t,t)(1-t,t) on the horizontal axis. The top subplot uses the function ot.gromov.gromov_barycenters from the POT Library [22] (based on [49]). The bottom subplot uses the blow-up technique from [14], which appropriately realigns the nodes and enlarges the template matrices $\mathbf{X}^{0}$ (zero shape) and $\mathbf{X}^{1}$ (one shape), creating new versions $\mathbf{X}^{0}_{b}$ and $\mathbf{X}^{1}_{b}$ of the same size, so that the GW barycenters can be interpreted as convex combinations $(1-t)\mathbf{X}^{0}_{b}+t\mathbf{X}^{1}_{b}$ (see Remark B.9 in Suppl. Mat. B.3). MDS embedding is used for visualization.
@@
From a machine learning perspective, a substantial literature has developed on the use of the GW distance for graph representation. In particular, the works [76, 75] introduce GW factorization models in which each graph is encoded as a barycentric combination of learned graph atoms, yielding permutation-invariant embeddings. Related extensions and applications can be found in [82]. In these approaches, the primary goal is to design neural network architectures that perform GW dictionary learning, so that the atoms (references or templates) are learned jointly with the associated weights. Consequently, the emphasis lies on the latent representation space and efficient learning algorithms. In supervised graph prediction [9], fused-GW barycenters models are used. A closely related line of work based on an online dictionary-learning framework [69] similarly trains a dictionary of GW and fused-GW atoms and embeds graphs in the corresponding atom-weight space by using a linear approach. Indeed, the model is a relaxation used for computational tractability: each data point $\mathbb{Y}$ is approximated as a convex combination of learned atoms with respect to GW distance ($\min_{\lambda\in\Delta_{S-1}}GW(\mathbb{Y},\sum_{s=1}^{S}\lambda_{s}\mathbb{X}_{s})$), where the authors explicitly impose the condition that the dictionary consists of pairwise relation matrices of the same size, but meanwhile, each dataset graph $\mathbb{Y}$ can have its own size. These frameworks consider an empirical form or, more precisely, a deep learning form of the analysis problem in the latent space, which is in contrast to our geometric and analytic approaches that pertain to the $\lambda$ rather than a latent representation thereof and in particular provide exact expressions for recovering $\lambda$ in certain cases. On the one hand, our framework is simpler in that the templates are fixed in advance; on the other hand, our aim is to characterize the analysis problem (barycentric weights recovery) through closed-form expressions. Our methodology is closer to that of [8], but with the GW geometry replacing the Wasserstein one. In their setting, the atoms (histograms) are fixed a priori, and the barycentric weights are obtained by solving OT-based optimization problems, rather than via learned encoders. Regression is then performed in the barycentric coordinate space through classical optimization, not by backpropagation through a deep model.
@@ 1.1. Summary of Contributions
    A gradient-based method motivated interpreting GW barycenters not necessarily as exact minimizers of (2), but rather as critical points of a functional on the GW space.
Regarding (1), we adopt the following simple idea, which can be applied beyond the context of GW. Consider a finite set of templates $\mathbb{X}^{s}=(\mathbf{X}^{s},\mathrm{p}^{s})$, $s=1,\dots,S$. Given a weight vector $\lambda\in\Delta_{S-1}$, consider GW synthesis problem (2), where the minimization is performed over $\mathbb{Y}=(\mathbf{Y},\mathrm{q})$, with fixed matrix size $M$ for $\mathbf{Y}$ and a fixed probability vector $\mathrm{q}$ [49]. Assume that the GW synthesis problem (2) can be solved via a fixed-point iteration of the form $\mathbf{Y}_{n}=\rho_{\lambda}(\mathbf{Y}_{n-1})$. Then, if $(\mathbf{Y},\mathrm{q})$ lies in the GW barycenter space generated by the templates, the matrix $\mathbf{Y}$ must be a fixed point of the corresponding iterative scheme. That is, there exists an unknown ‘coordinate’ vector $\lambda_{\mathbf{Y}}$ such that $\mathbf{Y}=\rho_{\lambda_{\mathbf{Y}}}(\mathbf{Y})$. To recover $\lambda_{\mathbf{Y}}$, we differentiate a chosen matrix norm $\|\mathbf{Y}-\rho_{\lambda}(\mathbf{Y})\|$ with respect to $\lambda$ and find its minimum at $\lambda=\lambda_{\mathbf{Y}}$. In summary, to generate GW barycenters, we employ a simplified version of the iterative scheme proposed in [49], and show in Theorem 3.3 that elements of the barycenter space are fixed points of this iteration. Building on this, Theorem 3.6 provides an exact recovery method for $\lambda_{\mathbf{Y}}$, for any $\mathbf{Y}$ in the GW barycenter space, formulated as a convex quadratic program and implemented in Algorithm 2. See Fig. 2 for an illustration of its applicability.
@@ 1.2. Organization of the Paper
Section 2 reviews the GW distance. Our main results, their proofs, and new algorithms are presented in Sections 3 and 4, which address the GW analysis problem via two approaches: a fixed-point method (Section 3) and a gradient-based method (Section 4). Section 5 presents experiments on 2D and 3D point clouds that serve as a proof-of-concept for the analytical results. Conclusions follow in Section 6. Supplementary material includes: additional context on the BCM framework in Section A; background proofs and examples in Suppl. Mat. B; computational complexity of the proposed GW analysis algorithms in Suppl. Mat. C; further experimental results in Suppl. Mat. D; experimental exploration beyond the GW-BCM hypothesis (i.e., for targets outside the GW barycenter space) E; and an additional discussion on the computation of GW and $W_{2}$ barycenters in Suppl. Mat. F. The code to reproduce the experiments of this paper is available online at https://github.com/RocioDM/GW_BCM.
@@ 1.3. Notation
In general, superscripts will be used for the templates and subscripts for the synthesized objects. We denote by $S]:=\{1,\dots,S\}$ the set of indices from $1$ to $S$. The space of weights $\Delta_{S-1}$, also known as a standard simplex or probability simplex, is $\Delta_{S-1}:=\{\lambda=(\lambda_{1},\dots,\lambda_{S})\in\mathbb{R}^{S}\mid\lambda_{s}\geq 0\,\,\,\,\forall s\in[S],\text{ and }\sum_{s\in[S]}\lambda_{s}=1\}.$ If $\mathrm{p}$ is a vector and $\mathbf{X}$ a matrix, when there is no ambiguity regarding whether we are referring to entries or coordinates, we will write $\mathrm{p}[i]=\mathrm{p}_{i}$ and $\mathbf{X}[i,j]=\mathbf{X}_{ij}$. We recall that the Frobenius norm of a real-valued square matrix $\mathbf{X}$ is defined as $\|\mathbf{X}\|_{\mathrm{Frob}}:=\sqrt{\mathrm{tr}(\mathbf{X}^{T}\mathbf{X})}$, where $\mathbf{X}^{T}$ stands for transpose. The component-wise (Hadamard) product between matrices is denoted by $(\mathbf{X}\odot\mathbf{Y})_{jl}:=\mathbf{X}_{jl}\mathbf{Y}_{jl}.$ For a vector $\mathrm{q}$, the matrix $\frac{1}{\mathrm{q}\mathrm{q}^{T}}$ is defined by $\left(\frac{1}{\mathrm{q}\mathrm{q}^{T}}\right)[{j,l}]:=\frac{1}{\mathrm{q}_{j}\mathrm{q}_{l}}.$
Given two (Borel) measure spaces $(X,\mu_{X})$ and $(Y,\mu_{Y})$, we say that the measure $\mu_{Y}$ is the pushforward of the measure $\mu_{X}$ by a measurable map $\varphi:X\to Y$, denoted by $\varphi_{\#}\mu_{X}=\mu_{Y}$, if $\mu_{Y}(A)=\mu_{X}(\{x\in X\mid\varphi(x)\in A\})$ for every measurable set $A\subseteq Y$. Given $\omega:Y\times Y\to\mathbb{R}$, we use the pullback notation $\varphi^{*}\omega(x,x^{\prime}):=\omega(\varphi(x),\varphi(x^{\prime}))$, for all $x,x^{\prime}\in X$. A coupling or transport plan $\pi$ between $\mu_{X}$ and $\mu_{Y}$, denoted $\pi\in\Pi(\mu,\nu)$, is a probability measure on the product space $X\times Y$ with marginals $\mu_{X}$, $\mu_{Y}$. In the case of finite discrete measures, $\mu_{X}=\sum_{i=1}^{N}\mathrm{p}_{i}\delta_{x_{i}}$ and $\mu_{Y}=\sum_{j=1}^{M}\mathrm{q}_{j}\delta_{y_{j}}$, each coupling can be identified with a matrix in $\Pi(\mathrm{p},\mathrm{q}):=\{\pi\in\mathbb{R}^{N\times M}\mid\,\pi_{ij}\geq 0,\,\forall i,j\,;\,\sum_{j\in[M]}{\pi_{ij}}=\mathrm{p}_{i}\,\forall i\,\text{ and }\sum_{i\in[N]}\pi_{ij}=\mathrm{q}_{j}\,\forall j\}.$